Paul is 7 years older than his brother. The product of their ages is 30. How old is Paul?
Answer by stanbon(75887)

Paul is 7 years older than his brother. The product of their ages is 30. How old is Paul?
---------------
Equations:'
p = b + 7
p*b = 30
----
Substitute for "p" and solve for "b":
(b+7)b = 30
b^2 + 7b = 30
b^2 + 7b - 30 = 0
Factor:
b^2 + 10b-3b - 30 = 0
-------
b(b+10) -3(b+10) = 0
(b+10)(b-3) = 0
Positive solution:
b = 3 yrs. (brother's age)
p = b+7 = 10 yrs (Paul's age)
=================
Cheers,
Stan H.
=================